# Social Media Isn’t Just Instagram: A Youth-Envisioned Platform for Meaningful Social Connections

### v1

@@ 3.1. Material: The Narrative Framework
Out of the numerous magical and fantastical worlds available, we identified the Harry Potter universe as particularly well-suited for our FI workshops. The series’ extensive popularity and deeply embedded cultural presence provide participants with a shared foundation of understanding, facilitating both recruitment and engagement. The Harry Potter series has garnered immense popular appeal and commercial success; by 2023, 98% of youth aged 18 to 27 in the United States identified as either casual or avid fans of the Harry Potter movie franchise, with approximately half describing themselves as avid fans (Share of Harry Potter fans U.S. 2023, [n. d.]). Furthermore, the Harry Potter book series has sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K. Rowling’s Harry Potter and the Sorcerer’s Stone, [n. d.]) and translated into over 80 languages by 2020 globally (The Rowling Library, 2020).
The series is set in modern-day England, reimagined to include magical elements, like enchanted objects and spells. Thus, this context leaves room for both incorporating or discarding any aspect of contemporary life, including designs and constraints related to social technologies. Its young protagonists and coming-of-age narrative naturally align with our study’s focus on youth identity and relationship building. The characters’ experiences or life transitions—which parallel real-world moves between schools or locations—allow for exploring how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging also lends itself to examining youth identity expression and self-presentation in virtual spaces.
@@ 3.2. Procedure
The first author conducted all the interviews over Zoom, with the second author joining one of the sessions to validate and ground the FI workshop method. The interviews averaged around 90 minutes. We began the interview by asking which 3D games they had played and what they liked about those experiences. We also asked questions about Harry Potter, such as their favorite book in the series. This served to break the ice and help them start immersing themselves in the Harry Potter world.
@@ 3.3. Participants and Recruitment
Participants were also required to be sufficiently familiar with the Harry Potter narrative such that they agreed with the statement, “I remember some details about the plot, characters, and magical aspects.” and to have prior experience with 3D games (e.g., Minecraft, Roblox). This last criterion was based on the assumption that individuals familiar with 3D games would likely have experience with 2D platforms as well, equipping them with a broader set of cognitive tools for creative idea generation. By contrast, extensive experience with 2D platforms alone does not necessarily ensure the ability to conceptualize or engage with spatial technologies effectively. While we did not impose constraints on social media experience, all participants had at least some engagement with social media.
@@ 3.4. Data Analysis
Our analysis began with the first two authors independently coding the same two transcripts line by line. During this phase, the codes were descriptive and closely adhered to the data. The authors convened to discuss and resolve any discrepancies in their coding. Following this, each author individually coded two different transcripts using Atlas.ti (atl, 2024). After four coding iterations, we reached saturation in our codes. Using this final set of codes, the first author proceeded to code all the transcripts.
@@ 4. Results
Participants articulated frustrations with current social media and envisioned Social Media at Hogwarts (SMH) as a transformative 3D platform prioritizing authentic connections with close friends. They proposed customizable personal spaces, communal areas fostering meaningful interactions, and immersive shared experiences like gaming or walking together. They suggested features that allow users to share expressively, such as memory sharing and real-time updates via avatars or aesthetics alongside low-intensity interactions that offer presence and nuanced privacy. Participants prioritized user control over feed content, fostering closer relationships over unsolicited interactions. Table 2 summarizes this section, while Figure 2 visualizes the collective vision, and Appendix  3 outlines key features per participant.
Figure 2. An AI-generated illustration envisioning a platform whose design reflects a synthesis of participants’ ideas. It is a 3D, open-world social media world with personal spaces—where users mirror their identity and interests—and community areas that support shared activities and memories. These homes exist in a neighborhood-based spatial design, where proximity mirrors social closeness, enabling intentional navigation and serendipitous interactions rather than passive scrolling. The neighborhood also includes themed social rooms and community spaces for public gatherings and interest-based group interactions. Users can visit friends’ homes and explore shared memories and content, similar to a Pensieve, with selective access control over different areas, rooms, and objects, ensuring privacy and customizable levels of sharing. An intelligent pet guide assists in exploration, connection, and engagement, fostering personal expression and meaningful interactions beyond traditional social media.
[…]
A whimsical digital illustration depicting a fantasy-style virtual neighborhood at night, with warm glowing houses, cobblestone pathways, and floating orbs of light. The left side shows a village scene, where small, cozy houses are arranged in a grid, surrounded by gardens, fences, and street lamps. A blue glowing node and arrow highlight one particular house, linking it to the right-side image. The right side provides a closer view inside a warmly lit home, where a young girl in a sweater and dress interacts with a cake on a wooden table. A small holographic memory scene emerges from the cake with a floating, glowing pool of light, depicting a group of people joyfully gathered around the cake. A curious orange dog watches nearby. Shelves filled with books, a spiral staircase, and a glowing lantern enhance the cozy, magical atmosphere. The image suggests an interactive, immersive social media experience, where users can explore virtual homes and relive shared memories.
Desired Feature Key Idea Example Quote
Immersive Experiences Through Spatial Elements (Section 4.1) “In 2D, you’re just observing… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” (P01)
[Section 4.1.1]
Bringing Realism and Presence Through 3D
Emphasized need for physical presence and embodied interactions beyond flat screens “You turn on your voice, chat… walking around a physical world… It feels like you’re actually in the area of these people; versus traditional social media like Instagram, there’s a separation of the screen.” (P06)
[Section 4.1.2]
Personal Spaces and Community-Centered Landscapes
Desired personal homes and neighborhoods that mirror real-world social structures “You each get your own house… anything you want.” (P03)
[Section 4.1.3]
Holographic Display and Virtual Reality
Sought more natural and immersive interfaces beyond traditional devices “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” (P01)
Organic and Intentional Social Interactions (Section 4.2) “The antonym to pressure, stress-free, community-oriented, wholesome, accessible, and accepting.” (P09)
[Section 4.2.1]
Flexible and Safe Spaces for Meaningful Conversations
Valued direct conversations in protected spaces “Actual conversations for real connections, individualized chatting… to get to know people… rather than just posting on a thread.” (P22)
[Section 4.2.2]
Community Gathering Places
Desired spaces for shared activities and casual encounters “a little card game area, shared experiences, something fun” (P06)
[Section 4.2.2] (1)
Activity Spaces for Shared Experiences
Valued collaborative activities and shared projects “Part of the fun is not only are we actually in the space, but doing something together… we’re experiencing something together, almost.” (P06)
[Section 4.2.2] (2)
(Public) Third Places for Casual Encounters
Desired spaces for serendipitous interactions “You strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” (P03)
[Section 4.2.3]
Ambient Co-Presence
Valued low-intensity connections through shared virtual spaces “At one point, literally, we sat for an hour in silence, listening to a whole playlist… we were just sitting there listening, vibing the music.” (P19)
[Section 4.2.4]
Effortful and Intentional Navigation
Saw physical movement as a meaningful way to engage “Ideally, SMH would have some way that you specifically have to search for stuff… where you have to intentionally go in looking for the account or the post versus it just being handed to you.” (P05)
Expressive and Lower-Stakes Sharing (Section 4.3) “I’m into multiple things—you could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” (P01)
[Section 4.3.1]
Expression Through Personal Area and Avatar Customization
Wanted customizable avatars and environments “decorat[ing] a house or customizing an avatar allows others to get to know more about my personality… through my style.” (P16)
[Section 4.3.2]
Real-Time Sharing: Capturing States and Changes
Desired ways to share real-time states “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that” (P04)
[Section 4.3.3]
Sharing Memories: Multisensory and Emotional Experiences
Wanted to share memories in an immersive and clearer way “jumping into a pool or taking the first bite of an ice cream sundae… from your brain, pull it out, and just be able to share that” (P04)
Granular, Intuitive, and Fun Privacy Mechanisms (Section 4.4)
“If people are posting more like their genuine real life and their experiences, privacy would be a little bit more complicated.”
[Section 4.4.1]
Space-Based Privacy
Wanted privacy boundaries based on spatial metaphors “A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate.” (P13)
[Section 4.4.2]
Playful Privacy
Desired engaging privacy mechanisms with elements of fantasy “Secret knock or an object that reveals a secret passage to… secret rooms.” (P11)
[Section 4.4.3]
Contextual Privacy
Wanted context-aware privacy settings “Each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” (P14)
[Section 4.4.4]
Invisibility
Sought robust blocking and visibility control “I don’t wanna be able to tell that the person I blocked is even around me.” (P22)
[Section 4.4.5]
Age-Appropriate Spaces
Emphasized need for age separation “People who are 50 should not be engaging with people who are 13.” (P21)
Refocusing Interaction Priorities for Meaningful Engagement (Section 4.5) “We need old Instagram back in 2010, when my mom was posting me and my brother just for the family and no one else. We didn’t need the explore page.” (P19)
[Section 4.5.1]
Intentional Content Consumption
Desired more control over content exposure “You press which floor you want to go to, and then, when you ride that elevator, it just goes straight to that floor.” (P09)
[Section 4.5.2]
Reduced Celebrity and Commercial Influence
Wanted focus on real friends over celebrities “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” (P07)
[Section 4.5.3]
Spatial Representation of Relationships
Sought physical distance to represent emotional closeness “The top 7 people you’re closest to would be the 7 houses in your cul-de-sac area.” (P15)
[Section 4.5.4]
Intentional and Meaningful Use of Time
Valued active engagement over passive consumption “Anytime not scrolling and instead talking to people feels worthwhile, even if it’s just light conversation.” (P06)
4.1. Immersive Experiences Through Spatial Elements
Participants saw 3D and VR platforms as key to emotionally engaging, socially fulfilling interactions. P16 noted current social media (which we will refer to as “Muggle Social Media” or MSM for short) “does a good job at spreading information [but not] emotions,” arguing that “tight interpersonal interactions” require “talk[ing] to you and see[ing] your face,” which 3D and VR enhance. They envisioned SMH as a 3D open world with realistic, highly flexible social interactions, lifelike avatars, distinct public and personal spaces, and holographic interfaces that convey presence.
4.1.1. Bringing Realism and Presence Through 3D
Many found MSM “flat” or limited by “just a screen,” (P10) yearning for greater realism and freedom. P19 compared SMH to “social media like Minecraft” but broader in scope, supporting a “physical world through 3D graphics” (P19) with “more of a connection” (P06) and “an emphasis on communications.” (P06) P06 referred to the 3D massively multiplayer online game (MMO) Sky (sky, [n. d.]), where players can “talk to other people” or “sit down by a candle and chat” but wanted “more of a connection” than “just sitting there, talking to one random person.”
Navigating shared virtual spaces resonated with P03, who noted that actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.” P06 similarly highlighted co-presence in 3D:
“it does feel more real… You turn on your voice, chat… walking around a physical world… It feels like you’re actually in the area of these people; versus traditional social media like Instagram, there’s a separation of the screen… you’re sharing the chat room. But you’re not really sharing the space [like when] I’m playing Roblox. [In SMH] I’m sharing that space physically, even though it’s not really physical.” (P06)
4.1.2. Personal Spaces and Community-Centered Landscapes
Participants wanted SMH’s 3D world to mirror the structure and functionality of real-world settings like homes, neighborhoods, and shared areas. P19 imagined “separate houses” with “a main house,” where users could “recreat[e] and decorat[e] your house.” P03 echoed this: “You each get your own house… anything you want” (P03). P21 expanded on the idea, seeing personal spaces as “subspaces” within a larger ecosystem.
Having a realistic landscape was essential for fostering connection and community for some. P06 described it as having “neighbors” or an “island next door” in a “town area.” This idea of community and proximity reflected participants’ desire to recreate the social structures of the real world in a digital setting. By mirroring the physical world, the landscape was also presumed to create a sense of place, grounding social interactions in a familiar and relatable context where users feel comfort and ownership. For example, P09 noted traversal in a corporal form would feel “fulfilling,” “calming,” and “reassuring.”
4.1.3. Holographic Display and Virtual Reality
A recurring theme was moving beyond traditional devices like phones and computers to embrace holographic and VR-based interfaces. Many participants felt mobile devices would seem cumbersome in a magical, immersive world. P01 noted, “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” They envisioned wands projecting holographic screens, like “Jarvis in Iron Man,” or using spell-activated “panels” and “big screen[s].” P15 preferred VR glasses, calling them “more natural,” “easier,” and “more fun.” Other concepts included a “makeup mirror” as a portal, a “watch” that immerses users, or a “mirror glass” for video calls or summoning friends via the charm (i.e., accio).
The immersive nature of VR was seen as fostering deeper connections. P01 noted MSM limits “how close you can be with someone,” explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.” They concluded 3D VR platforms are “more immersive and more awesome” (P03) than MSM, which is inherently limited “because it’s 2D.”
4.2. Organic and Intentional Social Interactions
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This carefully curated image culture creates “a lot of seriousness” (P23), which is reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared an anecdote about helping a friend prepare an Instagram post, including selecting the “post order,” highlighting “how seriously we take social media and our image online.”
In contrast, SMH was envisioned as a platform designed for intentional, engaging interactions with close friends, making it naturally less prone to issues like social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features and structures fostering organic, intentional, and personalized interactions, leading to closer friendships.
4.2.1. Flexible and Safe Spaces for Meaningful Conversations
Many participants valued direct, synchronous conversations as the best way to build closer relationships. P22 wanted, “actual conversations” for “real connections,” preferring “individualized chatting…to get to know people…rather than just posting on a thread.” They envisioned “chat rooms you can walk into,” (P02) each dedicated to specific topics, such as a room for “high school stuff,” decorated with artifacts from shared experiences to spark conversations. These were seen as “protected space…where you can talk…without feeling unsafe or judged” contrasting the perceived lack of safety on MSM, where users “don’t know everything around it.”
P09 imagined a flexible, “Room of Requirement” (roo, [n. d.])-like system for creating customized rooms “on a moment’s notice,” reflecting the “purpose” of conversations. P07 also imagined that personalized chatroom aesthetics could make conversations feel more intimate, such as “photos of you together, inside jokes, or fun memories.” Interest-based spaces, like Discord (discord page, [n. d.]), were also valued for providing a “common point” (P13) for conversation and “a good talking point for introductions.” (P07)
4.2.2. Community Gathering Places
SMH’s envisioned community spaces blended intimate, private interactions with dynamic public areas, fostering shared experiences and casual encounters.
Activity Spaces for Shared Experiences. Participants saw shared activities as central to friendship-building. P06 envisioned “a little card game area” like Roblox, where “shared experiences” of “something fun” create “real memories” to be reminisced, “do you remember that time when…?” They added, “Part of the fun is not only are we actually in the space, but doing something together… we’re experiencing something together, almost.” P12 valued collaborative “projects online,” such as “creat[ing] a house we love,” offering “a sense of self-achievement” and “teamwork” “instead of just like doom scrolling.” Sharing that experience together “on the same topic, on the same project” makes it “a better way to connect, instead of just… posting.”Even “mindless scrolling” became meaningful when shared. P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
(Public) Third Places for Casual Encounters. Although most of the interactions participants envisioned centered around friends, participants also emphasized public spaces for serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” Similarly, P08 suggested a marketplace space where selling items could spark conversations: “‘Hey, look! This is for sale’… have a talk about it… and meet new people.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you…and meet new people.” Other suggestions included “public study hall[s]” (P17) where participants could “interact with everybody,” (P17) leading them to experience that, “all of a sudden, you know this person that plays the piano.” (P17)
4.2.3. Ambient Co-Presence
Participants valued the idea of low-intensity connections through ambient co-presence. P18 described a “study sesh” where friends could send their avatars to “sit and study together” in a virtual study room with Lo-Fi music playing in the background. They explained that these spaces would mirror the 24/7 activity of Discord servers, offering a comforting presence even when not actively engaging in direct conversations. P06 highlighted the value of ambient interactions, emphasizing how these experiences allow people to “shar[e] that space physically, even though it’s not really physical.” Whether it was “experiencing the cafe together” or “read[ing] a book together” at a library, these shared moments of co-presence were described as uniquely meaningful: “We’re experiencing that together… just experiencing the space.”
Music would facilitate these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs.” They reflected on their experience:

### v2

@@ 3.1. Material: The Narrative Framework
Out of the numerous magical and fantastical worlds available, we identified the Harry Potter universe as particularly well-suited for our FI interviews. The series’ extensive popularity and deeply embedded cultural presence provide participants with a shared foundation of understanding, facilitating both recruitment and engagement. The Harry Potter series has garnered immense popular appeal and commercial success; by 2023, 98% of youth aged 18 to 27 in the United States identified as either casual or avid fans of the Harry Potter movie franchise, with approximately half describing themselves as avid fans (Share of Harry Potter fans U.S. 2023, [n. d.]). Furthermore, the Harry Potter book series has sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K. Rowling’s Harry Potter and the Sorcerer’s Stone, [n. d.]) and translated into over 80 languages by 2020 globally (The Rowling Library, 2020).
The series is set in modern-day England, reimagined to include magical elements, like enchanted objects and spells. Thus, this context leaves room for both incorporating or discarding any aspect of contemporary life, including designs and constraints related to social technologies. Its young protagonists and coming-of-age narrative naturally align with our study’s focus on youth identity and relationship building. The characters’ experiences or life transitions, which parallel real-world moves between schools or locations, allow for exploring how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging also lends itself to examining youth identity expression and self-presentation in virtual spaces.
@@ 3.2. Procedure
The first author conducted all the interviews over Zoom, with the second author joining one of the sessions to validate and ground the FI method. The interviews averaged around 90 minutes. We began the interview by asking which 3D games they had played and what they liked about those experiences. We also asked questions about Harry Potter, such as their favorite book in the series. This served to break the ice and help them start immersing themselves in the Harry Potter world.
@@ 3.3. Participants and Recruitment
They also had to be sufficiently familiar with the Harry Potter narrative such that they agreed with the statement, “I remember some details about the plot, characters, and magical aspects.” and to have prior experience with 3D games (e.g., Minecraft, Roblox). While we did not impose constraints on social media experience, all participants had at least some engagement with social media.
@@ 3.4. Data Analysis
Our analysis began with the first two authors independently coding the same two transcripts line by line. During this phase, the codes were descriptive and closely adhered to the data. The authors convened to discuss and resolve any discrepancies in their coding. Following this, each author individually coded two different transcripts using Atlas.ti (atl, 2024). After four coding iterations, we reached saturation in our codes. Using this final set of codes, the first author proceeded to code all the transcripts. All versions of the codes and the final codebook are available in the supplementary materials.
@@ 4. Results
Informed by their experiences with 3D games and familiar media worlds, our participants imagined forms of social interaction that extend beyond what is typically supported on mainstream platforms like Instagram. Through the speculative frame of Social Media at Hogwarts (SMH), they explored how digital environments might enable more embodied, relational, and emotionally meaningful forms of connection. Their visions converged around six interrelated themes, which, collectively, contribute to a vision of social media that is more meaningful, authentic, and aligned with their social and emotional needs, ultimately expanding their perception of the horizons for social connection on such platforms.
4.1. Intuitive Social Navigation
Youth described seeking interactions that feel natural, embodied, and socially meaningful, reflecting the way people navigate relationships and spaces in everyday life. They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape how they connect with others, express themselves, and manage privacy in intuitive and socially familiar ways.
4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
Participants described how embodied, immersive qualities such as realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space that fosters presence, emotional connection, and socially meaningful interaction. In contrast, current social media (which we will refer to as “Muggle Social Media” or MSM for short), was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 similarly noted that MSM limits “how close you can be with someone,” explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.” P03 described SMH as “more immersive and more awesome” than MSM, which they saw as inherently limited “because it’s 2D.” P16 emphasized that meaningful social connection requires “talk[ing] to you and see[ing] your face,” something they believed SMH could make possible.
For others, the fulfillment came from the embodied experience of physical actions. P04 valued “meandering through different people’s spaces,” finding that deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 echoed this, appreciating how moving physically between rooms created a “feeling of time passing.” They explained, “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They also added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
P03 similarly reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.” As P06 put it, SMH “feels more real” because you “walk around a physical world” and “feel like you’re actually in the area of these people.”
This sense of presence was felt even in seemingly passive actions. As P01 explained,
“I don’t know, there’s just a feeling about it that makes it different. I can’t really explain. Depending on 2D and 3D, but I feel like in 2D, you can’t really change anything about their homepage. You just observe it. The most you can do is add a like or something, and that would change their like count. But then they would know who you are, so that takes away the anonymity. But if you’re just looking at it, then you’re not really affecting anything. That’s just for your knowledge. But in a 3D setting, even if you don’t leave any physical footprints, in a way, they’re still there because you opened the door, you went in, you sat down. Maybe just that mark in a way that you’ve left on the place takes away from the anonymity. I don’t know if that makes sense but.” (P01)
For P01, 3D interaction not only supported more embodied engagement but also introduced a felt presence that lingered, even when no direct interaction took place.
4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
Participants proposed that using spatial and visual cues modeled after real-world environments could make managing access and privacy feel more intuitive, socially natural, and contextually appropriate. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
While managing visibility across multiple rooms might seem more complex, participants believed this approach would feel intuitive and natural. P11 suggested that spatial contexts could lower the efforts required for privacy controls, stating, “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
4.2. Shared Experiences
Participants emphasized that shared experiences, whether through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
4.2.1. Fostering Shared Memories Through Collaborative Activities
Participants described shared activities as a meaningful foundation for building friendships. P06 contrasted casual solo games with more immersive, co-experienced ones. For example, while playing a simple emoji game was described as “just something fun we can both do,” it lacked a true shared experience since “we don’t actually see each other’s screens” and would just “screenshot our scores and send them to each other.” By contrast, they described Roblox as a space that fosters shared moments that become lasting memories, like “do you remember that time when that kid in the voice chat said this,” or “when we got scared by that jump scare.” They explained, “it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
Similarly, P12 described the value of collaborative projects, such as “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting it with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than simply “posting” content individually.
The value of shared experiences and memories was not limited to completing tasks or reaching goals. Even simple activities like walking together were seen as meaningful opportunities for connection. P03 imagined moments where you could just “walk with your friend,” as if taking a “nature walk” (P14). P21 echoed this, describing how walking itself could spark memorable conversations and shared moments: “If I’m outside and we’re all in the mood to walk, then we can just choose to walk to the cafe… Because many times when you’re walking together, you come up with the most out-of-pocket conversations, and sometimes the best memories happen during a walk.”
4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
Participants also saw low-pressure, ambient forms of social interaction, such as co-existing in shared spaces or listening to music together, as natural and organic ways to build connections and shared memories. P18 described a “study sesh” where friends could send their avatars to “sit and study together” in a virtual study room with Lo-Fi music playing in the background. They explained that these spaces would mirror the 24/7 activity of Discord servers, offering a comforting presence even when not actively engaging in direct conversations. P06 highlighted the value of ambient interactions, emphasizing how these experiences allow people to “shar[e] that space physically, even though it’s not really physical.” Whether it was “experiencing the cafe together” or “read[ing] a book together” at a library, these shared moments of co-presence were described as uniquely meaningful: “We’re experiencing that together… just experiencing the space.” Even “mindless scrolling” became meaningful when shared. P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
Music would often facilitate these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs.” They came up with the idea reflecting on their own experience: